### v1

# On the Interchangeability of Positional Embeddings in Multilingual Neural Machine Translation Models

Abstract
Standard Neural Machine Translation (NMT) models have traditionally been trained with Sinusoidal Positional Embeddings (PEs), which are inadequate for capturing long-range dependencies and are inefficient for long-context or document-level translation. In contrast, state-of-the-art large language models (LLMs) employ relative PEs, demonstrating superior length generalization. This work explores the potential for efficiently switching the Positional Embeddings of pre-trained NMT models from absolute sinusoidal PEs to relative approaches such as RoPE and ALiBi. Our findings reveal that sinusoidal PEs can be effectively replaced with RoPE and ALiBi with negligible or no performance loss, achieved by fine-tuning on a small fraction of high-quality data. Additionally, models trained without Positional Embeddings (NoPE) are not a viable solution for Encoder-Decoder architectures, as they consistently under-perform compared to models utilizing any form of Positional Embedding. Furthermore, even a model trained from scratch with these relative PEs slightly under-performs a fine-tuned model, underscoring the efficiency and validity of our hypothesis.
On the Interchangeability of Positional Embeddings in
Multilingual Neural Machine Translation Models
Varun Gumma  Pranjal A. Chitale  Kalika Bali Microsoft Corporation
varun230999@gmail.com
1 Introduction
Neural Machine Translation (NMT) models have become integral to a broad spectrum of applications across various Natural Language Processing (NLP) tasks, particularly in generating synthetic data for training multilingual Large Language Models (LLMs). Despite their widespread use, state-of-the-art NMT systems often encounter sequence length constraints due to their reliance on Sinusoidal Positional Embeddings. Recent advancements have introduced alternative methods for embedding positional information, such as Rotary Positional Embeddings (RoPE) Su et al. (2024) and Attention with Linear Biases (ALiBi) Press et al. (2022), which demonstrate improved extrapolation capabilities for longer contexts. However, retraining existing models with these new embeddings is often computationally infeasible. This raises a critical question: Can we alter the Positional Embeddings in pre-trained models while preserving their performance? To the best of our knowledge, no prior study has examined the feasibility of swapping out Positional Embeddings in pre-trained transformer models.
This study addresses this question by investigating the feasibility of substituting Positional Embeddings in pre-trained NMT models. Additionally, inspired by prior works (Haviv et al., 2022; Kazemnejad et al., 2023), we explore whether Positional Embeddings are necessary in Encoder-Decoder models and whether these models can recover performance through fine-tuning after the removal of explicit positional information.
Figure 1: An basic overview of switching out Positional Embeddings
We conduct a thorough investigation into both parameter-efficient fine-tuning (PEFT) methods and full fine-tuning (FFT) strategies to restore performance and adapt pre-trained models to new Positional Embeddings. Our results demonstrate that it is indeed possible to substitute Positional Embeddings post-hoc without compromising model performance. Additionally, we find that adding extra parameters specifically to the self-attention block is sufficient for effective adaptation. This approach not only preserves sentence-level performance but also improves extrapolation capabilities for longer contexts when alternative Positional Embeddings, such as RoPE or ALiBi, are introduced.
In the remainder of this paper, we make the following contributions:
  • We provide a comprehensive analysis of the feasibility of modifying the Positional Embeddings of pre-trained transformer models post-hoc while maintaining comparable performance levels.
  • We demonstrate that these models can be adapted in a parameter-efficient manner by simply incorporating additional parameters into the self-attention block.
  • Our findings show that RoPE and ALiBi exhibit equivalent performance, while NoPE significantly degrades performance, highlighting it as a sub-optimal design choice for encoder-decoder transformer models.
  • We establish that post-hoc modification of Positional Embeddings achieves performance comparable to that of a model trained from scratch with the respective Positional Embeddings.
  • Lastly, we open-source111 https://github.com/VarunGumma/fairseq our framework to facilitate the efficient training and fine-tuning of long-context machine translation models.
2 Related Works
Multilingual Neural Machine Translation (MNMT)
Zhang et al. (2019); Firat et al. (2016); Aharoni et al. (2019); Arivazhagan et al. (2019) has emerged as a leading methodology for creating translation systems capable of managing multiple languages. These MNMT systems leverage a unified encoder-decoder transformer framework Vaswani et al. (2017) along with language-specific embeddings to integrate linguistic nuances. Previous works Fan et al. (2021); Costa-jussà et al. (2022) have developed many-to-many (X-Y) models at a global scale, while other works Ramesh et al. (2022); Gala et al. (2023) have primarily focused on many-to-one (X-En) and one-to-many (En-X) multilingual models for Indic languages. An in-depth examination of MNMT methodologies is presented in the survey by Dabre et al. (2020).
Positional Embedding Methods
Transformer-based models require an explicit Positional Embedding information as all the timesteps are processed parallely, and hence ther is a need for explicit positional. To this end, Vaswani et al. (2017) propose an empirical absolute Positional Embedding approach following a fourier series with alternating sine and cosine terms Kernes (2024). Devlin et al. (2019) introduce trainable Positional Embeddings which are learnt along with the other parameters of the model, however they have not shown a significant performance difference from prior static versions. Press et al. (2022) discuss ALiBi, a simple relative Positional Embedding strategy by adding the relative distance between tokens to the attention matrix and show that it can help the model generalize to longer sequence during inference even though trained on shorter sequences. RoPE Su et al. (2024) introduces the concept of rotating vectors according to their positions to explicitly encode relative position dependency in the self-attention formulation. Recent works Sun et al. (2023); Peng et al. (2023); Ding et al. (2024); Ostmeier et al. (2024) have also provided fixes to adapt RoPE to much longer sequences. Kazemnejad et al. (2023) find that no Positional Embedding can represents both relative and absolute strategies, and that it elicits high length generalizability. We refer the readers to Wang and Chen (2020) for a visual understanding of Positional Embeddings methods.
Parameter-Efficient Fine-tuning (PEFT)
Houlsby et al. (2019); Bapna and Firat (2019); Hu et al. (2022) deals with the concept of updating only a certain subset of parameters of the network, to efficiently generalize the model to new tasks or domains, instead of tuning all parameters, which might lead to catastrophic forgetting or over-fitting in large models. Low-Rank Adaptation (LoRA) Hu et al. (2022) is currently the most popular parameter efficient finetuning strategy, which hypothesizes that the update to weights are low-rank, and injects small trainable parameters directly into the pre-trained weights to mimic the low-rank decomposed updates. An extension, Weight-Decomposed Low-Rank Adaptation (DoRA) Liu et al. (2024) decomposes the pre-trained weight into a magnitude and direction, and employs LoRA only for directional updates to efficiently minimize the number of trainable parameters.
3 Methodology
In this section, we discuss the dataset curation, model training and evaluation procedures.
3.1 Model
Our experimentation focuses on standard Encoder-Decoder transformer models. While several massively multilingual models, such as NLLB (Costa-jussà et al., 2022) and its distilled variants, have been released, our experimentation requires the availability of long-context or document-level training and evaluation datasets. Consequently, we focus on language-family specific models that allow for more controlled experimentation. For our experiments, we chose the distilled variants of IndicTrans2 Gala et al. (2023). A key difference is that the distilled models released in Gala et al. (2023) we used were not trained with BackTranslation Sennrich et al. (2016) data (BPCC-BT), which we incorporate in our experimental setup. We experimented with models based on RoPE, ALiBi, and NoPE by replacing the sinusoidal Positional Embedding in the pre-trained model with the corresponding relative Positional Embedding. This replacement led to a significant performance drop, prompting us to fine-tune the models using high-quality data Mohiuddin et al. (2022); Gumma et al. (2023) to study their recovery in each setting.
3.2 Data
3.2.1 Train set
Fine-tuning
For maximum performance gains for our models, we choose a high-quality mixture of seen and unseen distributions for our training data. To this end, for the seen distribution, we select the BPCC-Seed, which is a high-quality human-verified subset BPCC. This subset consists of X pairs, and has shown to be effective in improving model performance Gala et al. (2023). As for the unseen distribution, we sample the top 150K pairs for each language from the BackTranslation (BPCC-BT) subset of BPCC based on the provided LASER Costa-jussà et al. (2022) or LaBSE Feng et al. (2022) cosine-similarity scores. In case similarity scores are unavailable for a translation pair, a random sample was chosen. Additionally, for the unseen data, we also include the Asian Language Treebank [^2] https://www2.nict.go.jp/astrec-att/member/mutiyama/ALT (ALT) Corpus Riza et al. (2016) which contains long-context translation pairs for Hindi and Bengali, and the CoPara Aligned corpus E et al. (2023), which contains document-level translation pairs for 4 Dravidian languages, Telugu, Kannada, Malayalam and Tamil. Overall, our train set consists of 4.45M and 3.46M translation pairs across 22 languages in the X-En and En-X directions respectively.
Training
We combine the fine-tuning data prepared above with the training data of the IT2 distilled models as described in Gala et al. (2023), to create 114.09M and 113.10M translation pairs for the X-En and En-X directions, respectively.
3.2.2 Dev/Test sets
Most existing benchmarks for Indian languages are composed of sentence-level test sets, which may not adequately reflect long-context performance or extrapolation capabilities. To address this, we construct document-level development and test sets using Flores [^3] https://huggingface.co/datasets/facebook/Flores (Costa-jussà et al., 2022) and IN22-Conv [^4] https://huggingface.co/datasets/ai4bharat/IN22-Conv (Gala et al., 2023). In this setup, sentences are grouped within the same context window, making them suitable for document-level evaluation. For Flores, we group sentences by metadata, including URL, domain, and topic headers, to map sentence-level pairs into document-level pairs, where each document-level unit consists of three sentences. This method is viable since this benchmark was created by professional translators who translate three consecutive sentences from every selected context window, as described by Goyal et al. (2022). In case of IN22-Conv, the dataset was originally created by translating entire conversations from English into the target languages by professional translators, and used for sentence-level evaluation by Gala et al. (2023). We use metadata provided to trace back the conversations and treat the entire conversation as a sequence for long-context evaluation. As a result, we obtain 289 merged sentence pairs per language, with a total of 6,380 pairs.
3.3 Metrics and Evaluation
For consistency, we use the ChrF++ score (Popović, 2017) as the primary metric, and follow the same evaluation practices as outlined in Gala et al. (2023) by normalizing the indic texts using indic_nlp_library [^5] https://github.com/anoopkunchukuttan/indic_nlp_library. For brevity, we report the aggregate score across all the languages in the respective test set, which are denoted in Table 1.
Test Set Languages Covered
IN22
Assamese, Bengali, Bodo, Dogri,
Konkani, Gujarati, Hindi, Kannada,
Kashmiri, Maithili, Malayalam, Marathi,
Manipuri, Nepali, Odia, Punjabi,
Sanskrit, Santali, Sindhi, Tamil,
Telugu, Urdu
Flores
Assamese, Bengali, Gujarati, Hindi,
Kannada, Kashmiri, Maithili, Malayalam,
Marathi, Manipuri, Nepali, Odia,
Punjabi, Sanskrit, Santali, Sindhi,
Tamil, Telugu, Urdu
ALT
Bengali, Hindi
Table 1: List of languages in each test set.
3.4 Finetuning Strategies
In our fine-tuning experiments, we swap the sinusoidal Positional Embedding module with a RoPE, ALiBi or NoPE module, and fine-tune the model as efficiently as possible from that point to analyse the performance recovery.
FFT
First, we choose a full-fine-tuning (FFT) approach which tweaks all the 211.78M parameters of the model. This would allow the embeddings, layer-normalizations, and the rest of the linear layers666self-attention, cross-attention, feed-forward modules to be tuned as per the new relative Positional Embedding.
LoRA
In vanilla PEFT method, we add trainable parameters only to the linear layers resulting in 6.48M (3.06%) trainable parameters. Note that, unlike FFT, in this case, the embeddings and layer-normalizations do not have any trainable parameters. The main hypothesis in this experiment is we intend to understand if the model can adapt to the change in the Positional Embeddings by addition of few trainable parameters as FFT may be computationally expensive as well as result in catastrophic forgetting.
minLoRA
The relative Positional Embedding modules are added to the self-attention modules at every layers. Therefore, we intend to leverage this property and understand whether, we can further cut down on parameters and apply these additional parameters only to the affected modules. Hence, we test the hypothesis of tuning only the self-attention modules with LoRA resulting in 2.35M (1.11%) trainable parameters.
4 Experiments
We use the fairseq [^7] Our variant implements, LoRA, RoPE, ALiBi, and a slightly efficient SDPA as per the requirements of this work framework Ott et al. (2019) for all of our experiments. All the finetuning experiments were conducted on 4 $\times$ A100 80Gb GPUs, the train-from-scratch experiments on 8 $\times$ H100 80Gb GPUs, and every model was trained till convergence with early-stopping. The exact hyperparameter sets are mentioned in Tables 2 and 3.
Hyperparameter Value
Global Max Tokens 32K
Optimizer Adam
LR scheduler inverse_sqrt
Gradient clip norm 1.0
Adam betas (0.9, 0.98) / (0.9, 0.999)
Checkpoint 1000 steps, BLEU @ $k=5$
Patience 10
Max Positions (4096, 4096)
Warmup steps 2000
Learning rate 6e-5 / 1e-4
Dropout 0.2 / 0
Label smoothing 0.1 / 0
LoRA rank 16
LoRA alpha 32
LoRA dropout 0.05
LoRA Rank Scaled True
FP16 True
Table 2: Default hyperparameter settings for finetuning. The cells with two numbers present the values for FFT/LoRA. LoRA finetuning acts on very specific modules, and adds minimal trainable parameters, while keeping the rest frozen. This acts like an implicit regularization, and hence we do not add any explicit strategies like label-smoothing and dropout to the model in this setup.
Hyperparameter Value
Global Max Tokens 256K
Optimizer Adam
LR scheduler inverse_sqrt
Gradient clip norm 1.0
Adam betas (0.9, 0.98)
Checkpoint 1000 steps, BLEU @ $k=5$
Patience 10
Max Positions (4096, 4096)
Warmup steps 4000
Learning rate 7e-4
Dropout 0.2
Label smoothing 0.1
FP16 True
Table 3: Default hyperparameter settings for training from scratch
4.1 Baselines
We compare our hypothesis against several baselines as outlined below. First, we examine a direct swapping approach, where models with sinusoidal Positional Embeddings are altered to not use Positional Embeddings (NoPE) or subject to alternate Positional Embeddings like RoPE, ALiBi, and their performance is evaluated. Additionally, we consider fine-tuning these models under three conditions outlined in section 3.4. Finally, to estimate a bound for each Positional Embedding method without swapping, we train transformer models from scratch using Sinusoidal, RoPE, ALiBi, and NoPE.
5 Results and Discussions
In this section, we discuss the results that aim to understand whether Positional Embeddings can be efficiently swapped posthoc in a trained model.
5.1 Swapping out Positional Embeddings
Tables 4 and 5 demonstrate the performance of several experimental setups that involve swapping of Positional Embeddings and followed by fine-tuning (if applicable).
Performance Degradation Post-Swap
As expected, a significant drop in performance is observed when Positional Embeddings are removed, compared to the based, with further deterioration occurring if the new type of Positional Embeddings are directly swapped in without subsequent fine-tuning. However, a key observation in this scenario is that performance does not completely collapse, suggesting that positional information may be partially encoded within the model itself. This indicates that attributing the entirety of positional information to the Positional Embeddings alone may not be entirely accurate.
Fine-tuning is key to performance recovery
When Positional Embeddings such as RoPE or ALiBi are introduced, across various fine-tuning setups, we consistently observe that the performance can be regained through fine-tuning. This demonstrates that it is feasible to modify Positional Embeddings at later stages and still retain the performance of the base model.
Efficacy of the Positional Embedding switch
A notable finding in this setting is that fine-tuning a model originally using sinusoidal embeddings yielded performance comparable to models fine-tuned after the switch to using RoPE or ALiBi embeddings. This further reinforces the conclusion that Positional Embeddings in pre-trained transformer models can be safely tweaked even in later stages. Although sinusoidal, rotary, and ALiBi embeddings show similar performance in long-context document-level tests used in the context of this study. The advantage of switching to RoPE or ALiBi is that further length extrapolation may now be feasible in these transformer models, as demonstrated by the findings of Su et al. (2024) and Press et al. (2022).
NoPE is not a solution in Encoder-Decoder models
Another key observation in our setting is that the findings of Kazemnejad et al. (2023) about explicit Positional Embeddings not being necessary does not generalize to Encoder-Decoder models. As we observe that no fine-tuning strategy is capable of bridging the gap and regaining performance.
Minimalistic Fine-tuning is sufficient
Across all the fine-tuning strategies evaluated, minimalistic LoRA fine-tuning, as well as FFT and LoRA, produced comparable performance. This suggests that extensive parameter adjustments are not necessary; adding LoRA parameters solely to the self-attention modules, where the relative Positional Embedding modules are integrated, is sufficient for adaptation. This approach is highly parameter-efficient, requiring only 2.35M trainable parameters (1.11%).
Base SINE NoPE RoPE ALiBi
IT2 N R A FFT L ML FFT L ML FFT L ML FFT L ML
Flores Indic-En 59.5 34.9 31.5 23.3 59.6 59.2 59.3 42.8 41.7 41 59.4 59.0 59.0 59.3 58.7 58.7
IN22-Conv Indic-En 55.8 38.6 38.0 35.0 55.8 55.6 55.6 44.8 44.3 43.7 55.6 55.3 55.4 55.3 55.1 55.2
IN22-Gen Indic-En 62.0 34.1 28.0 19.5 61.9 61.6 61.6 42.5 41.1 40.3 61.6 61.4 61.3 60.7 60.9 60.8
Flores En-Indic 47.9 31.5 27.0 15.9 48.1 47.7 47.5 39.0 37.6 36.6 48.2 47.7 47.5 49.4 47.6 47.5
IN22-Conv En-Indic 44.5 35.2 31.3 26.1 42.4 42.0 42.5 38.7 37.5 37.2 42.5 42.0 42.0 42.8 42.0 42.1
IN22-Gen En-Indic 48.3 30.1 24.9 13.4 46.3 46.1 46.7 36.9 35.1 33.7 46.3 46.2 46.0 46.4 46.0 45.9
Table 4: ChrF++ scores obtained when swapping out the Positional Embeddings and fine-tuning, evaluated across Sentence-level benchmarks. N denotes a direct transition (without any tuning) to No-Positional Embeddings, R denotes a direct transition to Rotary Positional Embeddings, and A denotes a direct transition to ALiBi Positional Embeddings. “FFT” represents Full Fine-Tuning, “L” signifies LoRA fine-tuning, and “ML” refers to minimalistic LoRA fine-tuning (applied exclusively to the Self-Attention block).
Base SINE NoPE RoPE ALiBi
IT2 N R A FFT L ML FFT L ML FFT L ML FFT L ML
Flores Indic-En 60.9 28.6 12.4 5.6 61.3 61.1 61.2 37.8 34.9 33.6 61.0 60.8 60.9 59.3 59.2 59
IN22-Conv Indic-En 27.5 10.2 2.3 1.1 32.7 32.0 31.5 15.1 13.4 13.7 33.0 32.8 33.0 30.2 29.7 28.1
ALT Indic-En 65.2 39.5 35.4 26.8 65.2 65.2 64.8 49.2 48.0 47.1 65.0 65.0 64.9 64.7 64.8 64.8
Flores En-Indic 49.5 23.8 14.8 4.8 49.8 49.2 49.1 33.1 29.2 26.8 49.9 49.1 49.1 49.4 48.6 48.3
IN22-Conv En-Indic 25.4 7.7 2.7 1.3 24.5 23.4 22.9 11.1 9.5 8.0 24.1 23.5 22.6 23.3 20.8 21.1
ALT En-Indic 54.9 36.7 31.9 19.4 55.3 55.0 54.9 45.6 43.9 43.1 55.3 55.0 55.0 55.4 55.0 55.0
Table 5: ChrF++ scores obtained when swapping out the Positional Embeddings and or fine-tuning, evaluated across Document-level benchmarks averaged across languages. N denotes a direct transition (without any tuning) to No-Positional Embeddings, R denotes a direct transition to Rotary Positional Embeddings, and A denotes a direct transition to ALiBi Positional Embeddings. “FFT” represents Full Fine-Tuning, “L” signifies LoRA fine-tuning, and “ML” refers to minimalistic LoRA fine-tuning (applied exclusively to the Self-Attention block).
NoPE RoPE ALiBi
Flores Indic-En 38.8 59.4 58.4
IN22-Conv Indic-En 14.0 33.2 31.6
ALT Indic-En 52.5 64.2 63.9
Flores En-Indic 37.2 49.3 49.1
IN22-Conv En-Indic 17.3 26.0 25.8
ALT En-Indic 48.4 55.2 55.2
Table 6: ChrF++ scores of models trained from scratch with various Positional Embeddings evaluated across Document-Level benchmarks averaged across languages.
NoPE RoPE ALiBi
Flores Indic-En 44.9 57.9 57.5
IN22-Conv Indic-En 45.5 54.2 54.0
IN22-Gen En-Indic 40.9 48.1 47.9
Flores En-Indic 40.6 47.6 47.6
IN22-Conv En-Indic 41.0 44.4 44.2
IN22-Gen Indic-En 44.9 60.3 59.9
Table 7: ChrF++ scores of models trained from scratch with various Positional Embeddings evaluated across Sentence-Level benchmarks averaged across languages.
5.2 Comparison with Training from scratch
Tables 6 and 7 demonstrate the performance trends of the models with the respective Positional Embedding trained from scratch. We compare these with the results in Tables 4 and 5. The comparison shows that models with swapped Positional Embeddings either outperform or are on par with their counterparts trained from scratch. However, training from scratch is resource-intensive, whereas the proposed approach of post-hoc Positional Embedding swapping proves to be a more efficient alternative in terms of both computation and data. Additionally, models trained from scratch using NoPE slightly outperform their swap counterparts, but remain suboptimal compared to any other Positional Embeddings (SINE / ALiBi / RoPE), further reinforcing the prior finding that NoPE is not a viable option for Encoder-Decoder models.
6 Conclusion
In this study, we have demonstrated that sinusoidal Positional Embeddings can be efficiently replaced with RoPE or ALiBi in a standard transformer model. We show that this substitution can be achieved in a parameter-efficient manner by simply introducing additional parameters within the self-attention module. Our findings also indicate that NoPE is not a viable solution for Encoder-Decoder models, underscoring the necessity of explicit positional information in such architectures. Although our experiments are focused on the machine translation (MT) task, we believe that this approach can be extended to other Encoder-Decoder models or Encoder-only models to enhance their long-context capabilities.
7 Limitations
Our work aligns with modular deep learning principles, exploring the post-hoc interchangeability of key components like Positional Embeddings. Although our preliminary findings suggest this is possible, several limitations should be noted:
  • All our experiments have been limited to a single model primarily due to the complete availability of relevant information regarding the IndicTrans2 model and dataset. Further the lack of availability of long-context data is also a limiting factor in performing this experimentation across languages.
  • Our investigation does not extend to examining the interchangeability between two relative PEs or from relative to sinusoidal PEs. This is due to the fact that older models predominantly employed sinusoidal PEs, whereas contemporary SOTA LLMs tend to favor RoPE, which has demonstrated robustness in handling long contexts. Therefore, a practically viable setup would involve a switch from sinusoidal to relative and not the other way around.
  • Our study concludes with demonstrating that interchangeability is feasible. Further, we also demonstrate that fine-tuning with sinusoidal embeddings also results in nearly similar performance. However, it is important to note that prior works demonstrate that RoPE provides superior length generalization capabilities, however as we do not have any long-context benchmark at our disposal as of now, we leave this aspect to future work.
  • Lastly, we also acknowledge that no specific hyperparameter tuning was conducted for this task, and we utilize the standard set available in the literature.
8 Acknowledgements
We thank Jay Gala (Research Assistant, MBZUAI), and Raj Dabre (Senior Researcher, NICT) for their invaluable insights and comments on paper. We also thank Amit Sharma (Principal Researcher, Microsoft Research India) for the discussions that helped us in refining the idea.